### v1

# PRISM-Loc: a Lightweight Long-range LiDAR Localization in Urban Environments with Topological Maps

Abstract
Localization in the environment is one of the crucial tasks of navigation of a mobile robot or a self-driving vehicle. For long-range routes, performing localization within a dense global lidar map in real time may be difficult, and the creation of such a map may require much memory. To this end, leveraging topological maps may be useful. In this work, we propose PRISM-Loc – a topological map-based approach for localization in large environments. The proposed approach leverages a twofold localization pipeline, which consists of global place recognition and estimation of the local pose inside the found location. For local pose estimation, we introduce an original lidar scan matching algorithm, which is based on 2D features and point-based optimization. We evaluate the proposed method on the ITLP-Campus dataset on a 3 km route, and compare it against the state-of-the-art metric map-based and place recognition-based competitors. The results of the experiments show that the proposed method outperforms its competitors both quality-wise and computationally-wise.
I INTRODUCTION
Localization is a fundamental requirement for mobile robots and self-driving vehicles, enabling them to accurately determine their position within complex environments. Traditionally, high-precision localization has relied on dense global LiDAR maps that provide detailed geometric representations of the surroundings. However, as the operational range increases, these dense maps become computationally expensive to process in real-time and require substantial memory resources. This challenge motivates the exploration of alternative representations, such as topological maps, which capture essential structural information with significantly lower memory demands.
Topological simultaneous localization and mapping (SLAM) methods represent the environment as a graph-based structure, where nodes correspond to key locations and edges define navigable paths between them. Unlike metric SLAM, which focuses on precise spatial coordinates, topological SLAM emphasizes structural and relational aspects of the environment. This approach enables more scalable localization while reducing computational and storage requirements.
Despite utilizing a graph-based structure, our approach still requires solving a rough pose estimation problem. This is achieved by matching incoming LiDAR scans against a database of precomputed topological nodes, each represented by a lightweight, image-like descriptor. This process yields a candidate location that approximates the vehicle’s global position.
In the second stage, we refine this estimation using a novel LiDAR scan matching algorithm. This algorithm, based on 2D feature extraction and point-based optimization, accurately determines the local pose within the identified candidate region.
In this study, we compare the proposed method against two state-of-the-art localization techniques using a large-scale outdoor urban dataset collected with the Clearpath Husky robot. The first method relies solely on point cloud data for localization, while the second integrates place recognition and scan matching to enhance accuracy and robustness. To evaluate the efficiency of these approaches, we analyze their computational and memory performance, comparing them with our proposed method. This assessment provides insights into the trade-offs between accuracy, efficiency, and scalability in large-scale urban environments, highlighting the advantages of our approach for real-world robotic applications.
To sum up, the contributions of this work are following:
  • Precise and lightweight localization method in topological maps which is based on place recognition, scan matching, and rough odometry estimation.
  • A novel lidar scan matching algorithm which estimates a precise relative transformation between two scans using the curbs detection.
  • An original algorithm of curbs detection based on raw lidar point clouds.
II RELATED WORK
II-A Long-range Mapping
The construction of large-scale metric maps for mobile robots has been extensively studied. One of the most common mapping methods is RTAB-Map [1], which employs a unique memory management technique for large-scale mapping. However, an evaluation conducted in [2] shows that, over long distances and with a noisy positioning source, RTAB-Map’s memory consumption increases significantly (up to 8 GB for a 1 km trajectory) and produces inconsistent maps with artifacts such as ”corridor bifurcation” (see Fig. 1).
Figure 1: RTAB-Map’s map with noisy positioning on a 1 km trajectory. One corridor is mapped twice.
To mitigate memory and computational costs, metric submap-based methods such as Voxgraph [3] can be utilized. The GLIM method [4] constructs a highly accurate metric map by representing the environment as a set of submaps, which are then stitched together into a voxelized point cloud. The authors tested this method on a 2 km trajectory in a park, achieving good trajectory estimation quality in real-time. However, memory consumption was not addressed in their study.
An alternative approach for long-range, large-scale mapping is the use of topological maps instead of dense global metric maps. This strategy has been implemented in works such as [5] and [6]. In [6], a topological map is represented as a graph of rooms, where doors serve as edges, and a metric map is constructed for each room separately. In [5], a city-scale topological map is built by fusing 2D global open-source maps with 3D point clouds obtained by a robot. This method can map a 17 $km^{2}$ urban indoor-outdoor environment while consuming only 1 GB of RAM. However, neither work provides a detailed description of the localization procedure within the topological map.
The TLF method [7] constructs a topological map based on multiple runs, facilitating lifelong localization under varying daytime, weather, and dynamic conditions. Localization is performed using a stereo camera, achieving an accuracy of 0.5 m and 4 degrees over multiple runs in different environments.
In recent years, learning-based topological mapping methods have emerged. For example, [8] introduces a framework for lifelong navigation, localization in a pre-built topological map, and map updates to accommodate dynamic changes. Localization is performed using a neural network that compares the current image with reference images from the topological graph. While such methods achieve high navigation success rates using stereo or RGB-D cameras, they are typically restricted to indoor environments and short trajectories.
II-B Place Recognition
To localize within a topological map, the system must first identify the correct location. This task is typically performed using place recognition techniques, which search for the location in the map whose observation best matches the current observation.
Early place recognition methods relied on the bag-of-words approach, where each observation is represented as an unordered set of descriptors. In visual place recognition, one of the most common bag-of-words-based methods is DBoW2 [9]. In the context of point cloud-based place recognition, BoW3D [10] has been successfully applied for loop closure on large outdoor trajectories.
More recently, learning-based place recognition methods have gained popularity. These methods encode an input image or point cloud as a single feature vector (descriptor) using a neural network. The correct location is then identified by comparing descriptor similarity between the current observation and map entries.
For vision-based place recognition, methods such as NetVLAD [11], CosPlace [12], and MixVPR [13] use convolutional or transformer-based architectures to predict descriptors. These methods achieve a Top-1 place recognition recall above 0.85 in outdoor urban environments.
For point cloud-based place recognition, PointNetVLAD [14] extends the NetVLAD architecture to 3D data. Other methods, such as MinkLoc3D [15] and its enhancement [16], utilize voxelization and sparse convolutions. SVT-Net [17] further refines this process by incorporating attention mechanisms. These approaches achieve recall rates of 0.87–0.93 in outdoor place recognition for urban environments.
Multimodal place recognition methods have also been developed, integrating image and point cloud data to improve performance. Examples include MinkLoc++ [18] and MSSPlace [19]. In our work, we utilize MinkLoc3D for place recognition as a lightweight and real-time method that relies solely on LiDAR point clouds.
II-C Pose Estimation
Once the correct location is identified through place recognition, pose estimation is performed using image or scan matching techniques. These methods extract features from the robot’s current observation and place recognition results, estimating the transformation between them.
For vision-based pose estimation, an example is StereoScan [20], which determines the transformation using stereo image pairs.
For point cloud-based pose estimation, a common approach is Iterative Closest Point (ICP) [21], which refines the transformation between two point clouds given an initial rough estimate. The RANSAC algorithm [22] is often used to generate this initial guess. However, the RANSAC + ICP pipeline is computationally expensive for dense point clouds and performs well only when the input point clouds have significant overlap. According to [23], this pipeline achieves a recall of 0.75 for point cloud pairs with overlap above 50%, but drops to 0.53 for pairs with overlap above 25%.
Using 2D projections of point clouds (e.g., bird’s-eye views) can improve pose estimation accuracy when overlap is low while also accelerating the matching process compared to full 3D point clouds. Methods such as BVMatch [24] and its enhancement BEVPlace++ [25] introduce a density-based bird’s-eye-view extraction algorithm and a novel local descriptor, achieving a pose estimation error of 0.4 m.
In our previous work [23], we introduced a fast bird’s-eye-view matching algorithm based on ORB feature matching with iterative outlier removal. In this work, we enhance our scan matching algorithm for outdoor environments by incorporating curb detection in point clouds and point-to-distance-map optimization.
III PROBLEM STATEMENT
Consider a mobile robot navigating in an outdoor urban environment using a pre-built map. During navigation, the robot does not have access to a global positioning source such as GNSS. Instead, it receives odometry information and point clouds from its LiDAR sensors at each time step. The objective is to estimate the robot’s position in the environment at each moment based on its sensor observations and the pre-built map. The robot’s initial position in the map is assumed to be known.
Formally, the localization problem is defined as follows:
During the first run, we collect input point clouds $C_{t}^{\text{first}}$ and the corresponding ground truth robot positions $p_{t}^{\text{first}}$ for each time step $t=1,\dots,T^{F}$. In subsequent runs, the robot receives new point clouds $C_{t}^{\text{next}}$ and odometry measurements $o_{t}^{\text{next}}$, which are subject to noise, for each time step $t=1,\dots,T^{N}$. The initial position $p_{0}^{\text{next}}$ is known, while the ground truth positions $p_{t}^{\text{next}}$ for $t=1,\dots,T^{N}$ remain unknown.
After the first run, we construct a map $M$ using the collected data:
$M=Mapping(C_{1}^{first},\dots,C_{T^{F}}^{first};p_{1}^{first},\dots,p_{T^{F}}^{first})$
The localization task at each step $t$ is as follows:
$\widehat{p_{t}^{next}}=Loc(M;C_{t}^{next};o_{t}^{next};\widehat{p_{t-1}^{next}})$
As the localization quality metric, we use the Absolte Trajectory Error (ATE). We compute mean and median error value over all the time steps:
$e_{t}=||\widehat{p_{t}^{next}}-p_{t}^{next}||_{2}$ (1)
$ATE_{mean}=\frac{1}{T^{N}}\sum\limits_{t=1}^{T^{N}}e_{t}$ (2)
$ATE_{median}=Median(e_{1},\dots,e_{T^{N}})$ (3)
Also we evaluate the localization success rate as the part of time steps where the localization error is lower than certain threshold:
$SR_{loc}=\frac{1}{T^{N}}\sum\limits_{t=1}^{T^{N}}I(e_{t}<Th)$ (4)
IV METHOD
IV-A Map creation
For map creation, we use the concept of a graph of locations and its implementation from PRISM-TopoMap method [23]. We build the map $M=(V,E)$, where $V=(v_{1},\dots,v_{N})$ are the locations with their ground truth poses $p_{1},\dots,p_{N}$ respectively, and $E=\{(v_{i}\in V,v_{j}\in V):\ ||p_{i}-p_{j}||<5\}$ are the edges between locations (we link each pair of locations whose observation points are located not further than 5 m from each other). As the relative poses for graph edges, we use the difference between ground truth poses of the locations. For each location, we store the 2D grid extracted from its point cloud by 2D projection, and the descriptor for place recognition. So, each location consumes only several kilobytes in disk storage, which lets us store even city-scale topological maps onboard. An example of a topological map is shown in Fig. 2.
Figure 2: An example of a topological map built by PRISM-TopoMap method (red nodes and blue edges), and a 2D grid of a location.
IV-B Localization
Figure 3: A scheme of the proposed global localization pipeline. It includes $F_{PR}$ place encoder (we use MinkLoc3D) and the proposed scan matching module. The output of the pipeline is the current location and the robot’s relative pose in this location.
IV-B1 Overall pipeline
The proposed localization pipeline maintains current state in the graph of locations. At each step $t$, this state is represented as the location $v_{cur}^{t}$ where the robot is currently located, and the robot’s relative position $T_{cur}^{t}$ in this location. The pipeline is similar to the graph maintaining pipeline of the original PRISM-TopoMap method [23], however, it has no part of a new node addition and checks the localization results for proximity to the current state:
  • First, we check whether the robot is still inside $v_{cur}^{t-1}$ and its current scan overlaps with $v_{cur}^{t-1}$ by a sufficient percent. If the check passes, we apply the odometry measurement to $T_{cur}$: $v_{cur}^{t}=v_{cur}^{t-1};T_{cur}^{t}=T_{cur}^{t-1}\cdot o_{t}.$
  • If we moved outside $v_{cur}^{t-1}$, or our overlapping percent is low, we first try to change $v_{cur}$ to one of its neighbors in the graph (i.e. move along an edge). For this purpose, we use the approximate relative position $((T_{cur}^{t-1})^{-1}(p_{v_{cur}})^{-1}p_{v_{next}})^{-1}$ as an initial guess for scan matching. If we matched successfully to one the neighbors $v_{next}$, then we change $v_{cur}$ to it: $v_{cur}^{t}=v_{next};T_{cur}^{t}=T_{next}$. To prevent false matches, we use a ”jumping threshold” – the maximum distance between an initial guess and the found transform when the matching is considered successful.
  • If we could not move along an edge, we try to change $v_{cur}$ using the localization results. If there is found the location $v_{loc}$ with sufficient overlap, we change $v_{cur}$ to it: $v_{cur}^{t}=v_{loc};T_{cur}^{t}=T_{loc}$.
  • After all, if we could not change $v_{cur}$ by edges and by the localization results, we change its value to the nearest neighbor in the graph without transform alignment (i.e. try to move along an edge without matching): $v_{cur}^{t}=v_{next};T_{cur}^{t}=((T_{cur}^{t-1})^{-1}(p_{v_{cur}})^{-1}p_{v_{next}})^{-1}$. We do this change if $||((T_{cur}^{t-1})^{-1}(p_{v_{cur}})^{-1}p_{v_{next}})^{-1}||<5$ to stay inside the next location. Otherwise, we do not change $v_{cur}$, apply step (1), and consider that the localization is lost.
The estimated global position in each time step is simply the application $T_{cur}^{t}$ to $v_{cur}^{t}$’s pose:
$\widehat{p_{t}^{next}}=p_{v_{cur}^{t}}^{first}T_{cur}^{t}$
For global localization in the topological map, we use a twofold pipeline which first searches the proper location in the map using place recognition, and then refines the place recognition results and provides the precise pose estimation using the proposed scan matching algorithm. The scheme of the pipeline is depicted in Fig. 3.
IV-B2 Place Recognition
For place recognition, we use MinkLoc3D model [15]. For each scan from the robot, we find top-5 nearest locations in the topological map using proximity of the descriptors extracted by the MinkLoc3D model. We then identify the top-5 most similar place descriptors within the nodes of the location graph based on Euclidean distance. These five candidates are utilized in the subsequent stages of our pipeline.
IV-B3 Scan Matching
Learning-based place recognition methods sometimes output false positive matches due to visual similarity of different parts of the environment. Therefore, relying solely on a place recognition network can cause our topological mapping method to link far, non-adjacent locations. To prevent this, we filter place recognition results matching the scan from the robot with the scans of the found locations, thereby finding relative transforms between these scans.
To estimate the transform between the reference scan $R$ from the robot, and candidate scan $C$ from the map, we combine three techniques. First, we match the scans using the features $R_{f}$ and $C_{f}$ extracted by ORB detector [26] (for global localization) and by Harris corner feature detector [27] for matching along an edge. We find the correspondences between features using FLANN matcher [28], as in the PRISM-TopoMap method.
Also, we create distance maps ${}^{c}D,^{w}D$ from curb and wall points of the candidate scan. Within these distance maps, we optimize the distances from the curbs and walls points of the current robot’s scan $(^{c}P,^{w}P)$ to the nearest curbs and walls on the candidate scan. To do this, we perform weighted Gauss-Newton optimization of the function consisting of three components:
  1. 1.
    $PointToPointFactor(\Omega_{f},R_{f},C_{f})$) is the square distance between the feature correspondences;
  2. 2.
    $PointToDistMapFactor(\Omega_{w},^{w}P,^{w}D)$ is the square distance from each wall point of the reference scan to the nearest one in the candidate scan;
  3. 3.
    $PointToDistMapFactor(\Omega_{c},^{c}P,^{c}D)$ is the square distance from each curb point of the reference scan to the nearest one in the candidate scan;
Figure 4: Scan matching comparison: our previous algorithm (left) [23] and the proposed algorithm (right). The candidate scan is shown in gray and white, the reference scan is shown in green (wall points) and red (curb points)
To make our scan matching more robust, we optimize iteratively, removing outliers from feature correspondences. We use a ”funnel” of thresholds to filter out outliers. The procedure of the matching is described in the Algorithm 1. An example of the comparison between the proposed scan matching and the PRISM-TopoMap’s one is shown in Fig. 4.
1
Data: $R_{f}$, $C_{f}$, ${}^{w}D$, ${}^{c}D$, $\nabla^{w}D$, $\nabla^{c}D$, ${}^{w}P$, ${}^{c}P$
Result: $X\in SE(2)$ is transform $T_{C,R}$, where C and R are frames attached to candidate grid and reference grid respectively
2 $X\leftarrow\text{transform from points $R_{p}$ , $C_{p}$}$ ;
3 for $\delta\in(15,7,1,0.1)$ do
4
5      $X\leftarrow$ Optimize ($X$, [
6       PointToPointFactor($\Omega_{f}$, $R_{f}$, $C_{f}$),
7       PointToDistMapFactor($\Omega_{w}$, ${}^{w}P$, ${}^{w}D$, …),
8       PointToDistMapFactor($\Omega_{c}$, ${}^{c}P$, ${}^{c}D$, …)
9       ], $N_{iter}$ = 10) ;
10
      // Outliers filtering
11       remove pairs of point $(r,c)$ $r\in R_{p}$ $c\in C_{p}$ from $R_{p}$ $C_{p}$ if $||\tau(X,r)-c||_{2}>\delta$ ;
12
13
// optimization use only walls, if we did not match
14 if $IOU(\tau(X,R),C)<0.4$ then
15
16      $X\leftarrow$ Optimize ($X$, [
17
18      PointToDistMapFactor($\Omega_{w}$, ${}^{w}P$, ${}^{w}D$, …)
19       ], $N_{iter}$ = 20) ;
20
Algorithm 1 Scan matching optimization
Figure 5: Curb detection output. Points belonging to curbs are shown in red.
IV-B4 Curb Detection
The key idea of the curb detection algorithm is to accurately segment the ground (drivable region) through plane fitting, from which curbs are derived as edges of this region. The process begins by applying a sliding window across the azimuth to compute the standard deviation of the radial distance (range) for each point. Points exhibiting a low standard deviation are retained, as they are likely to belong to flat surfaces such as the ground. Subsequently, points sampled from the previous step are weighted with a higher probability near the center line to minimize the influence of pavements on the sides. Plane fitting is then performed on these filtered points using the RANSAC algorithm. The original point cloud is rotated to align the vertical axis with the normal of the fitted plane. The ground region is defined as points lying below a threshold of 0.1 m and is represented as a binary 2D array (mask) in the range view. Edges within the ground region mask are detected while ignoring points near regions with heights exceeding 0.3 m to eliminate noise from dynamic objects such as pedestrians. Finally, the resulting indices of edges in the range view are used to reconstruct the 3D coordinates of the detected curbs in the robot‘s frame. Examples of curb detection algorithm output is shown in Fig. 5.
Data: Point cloud: $P$
Result: Points of curbs: $C$
// Ground plane segmentation
1 $P_{f}\leftarrow rangeSTDWindowFilter(P,\sigma_{threshold}^{2})$;
2 $P_{s}\leftarrow sample(P_{f},\sigma_{y}^{2})$;
3 $T_{pb}\leftarrow ransacFitPlane(P_{s},d_{threshold})$;
4 $P_{t}\leftarrow transform(P,T_{pb})$;
// Edges detection
5 $P_{u}\leftarrow upsample(P_{t},f=3)$;
// Mask in range view
[…]
8 $mask_{2}\leftarrow dilate(mask_{2},ones(w=50,h=7))$;
9 $mask\leftarrow\neg mask_{2}\wedge mask_{1}$;
$C\leftarrow edges(P_{u},mask)$
Algorithm 2 Curb Detection
V EXPERIMENTS
V-A Setup and Dataset
We evaluated the proposed method and the competitors on the outdoor part of the ITLP-Campus dataset [29]. This dataset contains several runs of a mobile differential-drive Husky robot on a university campus. The length of each run varied from 3 to 4 km and from 1 to 1.5 hours. These runs were recorded in different seasons and daytimes, and all of them cover the whole campus.
For all the methods, we used track 03 for map creation, and tracks 00, 01, and 02 for evaluation of the localization within the built map. The most challenging are the tracks 00 and 01 which were recorded in winter (whereas track 03 was recorded in spring). In addition to the dynamic changes between the tracks (parked and moving cars, pedestrians, etc.), the tracks 00 and 01 contain snow piles with heights up to 1 m.
V-B Baselines
We used 2 methods to compare with the proposed method.
The first one is a classical point cloud based approach, where the algorithm operates with a point cloud itself. For localization in such methods, an accurate and dense point cloud should be given as an input. GLIM [30] was used for creating a 3D map. Unfortunately, we cannot use GLIM localization, because it is a proprietary solution. Instead of it, we used HDL_Localization method [31] to localize within the GLIM-built map.
GLIM method uses fixed‐lag smoothing combined with a keyframe‐based point cloud matching approach to robustly estimate sensor trajectories. In addition, multi‐camera visual constraints are incorporated to further enhance stability and accuracy. The authors used the GPU-accelerated scan matching process, which enables the system to directly minimize registration errors between submaps in a densely connected factor graph – circumventing some of the limitations of conventional Gaussian pose graph optimization – while still achieving real-time performance despite its high computational load.
Another method was BEVPlace++ [25]. It introduces a fast, robust, and lightweight LiDAR global localization method designed for unmanned ground vehicles. The approach begins by projecting sparse LiDAR point clouds into bird’s-eye view (BEV) images, which serve as a compact and informative representation. A specially designed Rotation Equivariant and Invariant Network (REIN) then extracts rotation-equivariant local features and rotation-invariant global descriptors from these BEV images. This enables a two-stage localization process where coarse place recognition is performed via global descriptors and refined through local feature matching with RANSAC for accurate 3-DoF pose estimation. The method’s efficiency, combined with its robustness to view and sensor variations without requiring dense pose supervision, makes it an appealing baseline for real-time global localization tasks.
A significant limitation of the BEVPlace++ algorithm stems from its dependence on an approximate initial pose estimate. Specifically, the method requires prior knowledge of the vehicle’s rough global position at each timestamp to retrieve the optimal BEV reference image from the precomputed database. While this approach is effective for refining localization in systems equipped with supplementary sensors (e.g., wheel odometry), it inherently restricts the method’s applicability to global localization tasks using exclusively LiDAR-derived data. As illustrated in Figure 6, unfiltered BEVPlace++ exhibits an average Absolute Trajectory Error (ATE) Root Mean Square Error (RMSE) of approximately 73 meters across test trajectories, underscoring the necessity for robust filtering mechanisms.
(a) Trajectory on $XY$ plane
(b) Coordinates respect to time
Figure 6: BEVPlace++ method without applied filters, track 02
To overcome such limitations, two filtering strategies were devised, both leveraging the ”last known database position index” – a concept predicated on the assumption that sequential database indices correspond to roughly proximate poses. This approach circumvents reliance on external localization modalities (e.g., odometry) by exploiting temporal continuity and the known approximate inter-frame distance between consecutive database entries.
The first filter (denoted as ”BEVPlace++” in results) enforces temporal coherence by constraining the search space to database indices incrementally adjacent to the last confirmed position. Candidate poses retrieved via FAISS similarity search are evaluated based on their index proximity to the prior validated index. While this method improves robustness against isolated outliers, it does not account for loop closure events, leading to potential track loss.
TABLE I: Results on ITLP-Campus dataset
Method Mean ATE, m Median ATE, m $SR_{loc}$ RAM usage, GB Map size, GB Localization time, s
GLIM + HDL loc. 3.6* 3.5* 0.3 11 + 1.2** 0.4 3.0
BEVPlace++ 14.4 11.0 0.84 15.5 15.3 1.0
BEVPlace++ (outlier filtering) 20.3 6.8 0.95 15.5 15.3 1.0
PRISM-Loc 8.2 0.7 0.98 4.1 0.2 0.5
* ATE was estimated only before the localization failure
** 11 GB of RAM was used by GLIM for map creation, and 1.2 GB of RAM was used by HDL for localization
Filter 2 (denoted as ”BEVPlace++ wo outliers) incorporates spatial distance thresholds between the current pose estimate and the last validated pose. To mitigate cumulative drift caused by consecutive outliers, a buffer of recent outlier indices and their associated positional offsets is maintained. When a given distance threshold is exceeded, the system records the indices and distances of the outliers. If number of consecutive outliers have been scored, the filter pulls out the index with the minimum distance and makes it the last index. However, this method requires careful calibration of two hyperparameters: the size of the emission buffer and the threshold distance, which directly affect the robustness to stable emission sequences.
Two baseline frameworks were evaluated against the proposed method: (1) HDL localization based on GLIM created map, and (2) BEVPlace++ with 2 outlier filtering algorithms, which were developed for rough real-time pose estimation.
V-C Results
We recorded the positions estimated by each of the methods and measured the metrics 2 - 4 for them, as well as performance values, such as the map size in the disk storage, memory usage during the localization, and the average time spent to one global localization iteration. For computation of localization success rate, we use the threshold of 10 m.
The results are shown in Table I. The comparison of the estimated trajectories with all the methods and the boxplots for trajectory errors are shown in Fig. 7. At each track, HDL method failed to localize in the GLIM’s map after traveling about 1 km. The sparse voxelized point cloud map for the 3 km route weighed only about 400 MB; however, each localization procedure in it took about 3 s, with complete localization loss in the middle of the track. Because of significant changes between the tracks, the pure BEVPlace++ method without filtering exhibited about 16% of place recognition errors, which led to a mean ATE of 20 m (see Fig. 6). The suggested filtering approach eliminated most of the place recognition errors, however, the percent of localization fails remained high. The map for BEVPlace++ took 15 GB on disk storage and in the RAM because of storing large local descriptors for each location. Also, the localization pipeline took in average 1.0 sec (about 700 ms for place recognition, and 300 ms for pose estimation using a local descriptor).
The proposed PRISM-Loc method made localization losses and errors in only 2% of cases thanks to our combined approach to localization. The median ATE reached the value of 0.7 m, significantly outperforming all the competitors. With such localization quality, our approach remained fast and lightweight – the map of all the campus consumed only 200 MB (the largest part of 4 GB RAM consumption is the reserving allocation for PyTorch models, this memory is not utilized, so memory usage could be optimized in the future). Also, the global localization time took only 0.5 s (0.1 s for place recognition, and 0.08 s for scan matching with each of five found candidates).
(a) Trajectory on $XY$ plane
(b) Box Plot
Figure 7: Comparison analysis for track 00
VI CONCLUSION AND FUTURE WORK
We propose PRISM-Loc – a lightweight and robust method for localization in large outdoor environments using topological maps. The proposed method combines tracking through rough odometry estimations, a novel robust scan-matching technique integrated with the original curb detection algorithm, and place recognition for global location search. Experiments conducted on a challenging multi-season dataset covering a 3 km route demonstrate that the proposed method achieves robust localization, with a median trajectory error of 0.7 meters, despite significant object-wise and seasonal differences between the run and the map. Furthermore, PRISM-Loc maintains low memory usage and computational time throughout the entire route. Moving forward, we plan to enhance our approach by incorporating local pose filtering to improve robustness and to conduct extensive research on scan-matching techniques.